Circular No. 6931

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J2012+381 A. Castro-Tirado, Laboratorio de Astrofísica Espacial y Física Funda-

mental (LAEFF-INTA), Madrid, and Instituto de Astrofísica de Andalucía, Granada; and J. Gorosabel, LAEFF-INTA, report: “Following the detection of the x-ray transient XTE J2012+381 (IAUC 6920), U- and R-band images were taken on May 28.17 UT at the 2.2-m telescope at the German-Spanish Calar Alto Observatory. The images included the ASCA position reported on IAUC 6927. New U- and R-band images were obtained on June 3.13 by N. Walton (Isaac Newton Group of Telescopes, La Palma) at the 1.0-m Jacobus Kapteyn telescope. No variable (by > 0.2 mag) source is seen when comparing both images, down to a limiting magnitude of R = 21.5. We note that a star of mag 18 is coincident (within 1′′) with the variable radio source reported on IAUC 6924 and 6926.”

M. R. Garcia, J. E. McClintock, and E. Barton, Harvard-Smithsonian Center for Astrophysics; and P. Callanan, University College, Cork, report: “On June 1.45 UT, we obtained an optical spectrum of the USNO A1.0 star 1275.13846761, which is within ∼ 1′′.3 of the variable radio source (IAUC 6924, 6926), and in turn may be related to the x-ray transient XTE J2012+381 (IAUC 6920, 6922). Preliminary reduction of the 20-min exposure with the Whipple Observatory 1.5-m telescope (+ FAST spectrograph; 0.3-nm resolution, range 400–700 nm) shows a nearly featureless red continuum. The only clearly identifiable lines are Hβ, Hγ, and Na D in absorption; a strong night sky line at Hα makes it difficult to determine the intrinsic Hα profile. The upper limit to the equivalent width of any possible He II 468.4-nm emission is ∼ 0.1 nm. A cross-correlation with the Jacoby stellar atlas indicates a spectral type of F3 ± 3 subclasses for a dwarf, or F0 ± 6 subclasses for a giant. CCD photometry with the Whipple Observatory 1.2-m telescope on May 28.31 yields B = 19.35 ± 0.1, V = 18.1 ± 0.1, indicating B − V = +1.25. The photometry and spectroscopy are consistent with an absorption A(V) = 2.8, and an F3 dwarf at ∼ 3 kpc or an F0 giant at ∼ 6 kpc. These results neither confirm nor refute the possible association of this star with the x-ray nova. If this star is the counterpart, it could be similar to the F-giant companion of Cyg X-2, which shows weak (equivalent width ∼ 0 to 0.3 nm) and variable He II 468.6-nm emission, and sometimes an absence of H-Balmer emission. Alternatively, the star could be a chance coincidence along the line of sight, similar to the case of GX 17+2.”

1998 June 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green